OpenStudy (anonymous):

Find the point-slope form of the equation of the line satisfying the given conditions and use this to write the slope-intercept form of the equation. Slope = 6, passing through (2, 2)

OpenStudy (anonymous):

Point slope form : y – y1 = m(x – x1) m( is your slope) and x1 and y1 r your coordinates

OpenStudy (anonymous):

so what is the equation of the line?

OpenStudy (anonymous):

So then you get y-2=6(x-2) the equation n point-slope form now convert to slope intercept form y= mx+b :)

OpenStudy (anonymous):

y-2=6(x-2) point slope form Solve y-2 = 6x-12 Add 2 to both sides y= 6x-10 This is your equation in slope intercept form :)

OpenStudy (anonymous):

@bobiesco do u understand ?

OpenStudy (anonymous):

@bobiesco any q's?? :D